# Heterochromatic Geometric Transversals of Convex sets

### v4

@@ Abstract
An infinite sequence of sets $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence from an infinite sequence of families $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$, if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. In this paper, we have proved that if for each $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is a family of nicely shaped convex sets in $\mathbb{R}^{d}$ such that each heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ contains at least $k+2$ sets that can be pierced by a single $k$-flat ($k$-dimensional affine space) then all but finitely many families in $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ can be pierced by finitely many $k$-flats. This result can be considered as a countably colorful generalization of the $(\aleph_{0},k+2)$-theorem proved by Keller and Perles (Symposium on Computational Geometry 2022). We have also established the tightness of our result by proving a number of no-go theorems.
@@ 1 Introduction
Danzer, Grünbaum and Klee [DGK63] showed that the "congruent" assumption in Theorem 1.2 can be weakened as long as the sets are of bounded diameter, that is, there exists $R>0$ such that the diameter of all the convex sets in the family is at most $R$.
@@
Observe that for $k$-transversal $q(k)=k+2$ is optimal since any $k+1$ nonempty subset of $\mathbb{R}^{d}$ can be pierced by a single $k$-flat. Keller and Perles proved a similar result for point transversal of families of closed balls (with no restriction on radii). They also gave an explicit example showing that the compactness assumption in Theorem 1.7 cannot be weakened. Additionally, Keller and Perles observed that, unlike $(p,q)$-property, for all $10\leq k\leq d-$ 1 there does not exist a constant $c(d,k)>0$ such that for all families of unit balls in $\mathbb{R}^{d}$ satisfying $(\aleph_{0},K+2)$-property with respect to $k$-flats then these families have a $k$-transversal of size at most $c(d,k)$.
Given $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ families of convex sets in $\mathbb{R}^{d}$, we call a collection $B_{1},\dots,B_{p}$ of $p$ convex sets colorful with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ if for all $i\in\left\{1,\dots,p\right\}$ we have $B_{i}\in\mathcal{F}_{i}$. We say $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ satisfies colorful $(p,q)$-property with respect to $k$-flats if every colorful collection of $p$ sets with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ contains $q$ sets that can be pierced by a single $k$-flat. Lovász observed the following interesting variant of Helly’s theorem, called the Colorful Helly’s theorem, in 1974, though the proof first appeared in a paper on Carathéodory’s theorem by Bárány [B8́2].
Theorem 1.8 (Bárány [B8́2]: Colorful Helly’s theorem).
@@
Bárány, Fodor, Montejano, Oliveros, and Pór [${}^{+}$14] proved a colorful $(p,q)$-theorem for families of convex sets in $\mathbb{R}^{d}$.
Theorem 1.9 (Bárány-Fodor-Montejano-Oliveros-Pór [${}^{+}$14]: Colorful $(p,q)$-theorem).
@@
In contrast with Theorem 1.9, we shall show that no colorful $(\aleph_{0},q)$-theorem exists for $k$-flats and families of unit balls, except for the case where $k=0$. However, by expanding the definition of ‘colorful’, we can obtain a weak ‘colorful’$(\aleph_{0},k+2)$-theorem for $d>k>0$ involving families of unit balls and $k$-flats. For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of compact convex sets in $\mathbb{R}^{d}$. An infinite sequence of compact convex sets $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. A heterochromatic sequence is said to be a strictly heterochromatic sequence if each family has a representative in the sequence, that is, a heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a strictly heterochromatic sequence if for each $n\in\mathbb{N},\;B_{n}\in\mathcal{F}_{n}$.
@@
A sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets in $\mathbb{R}^{d}$ is said to be $k$-dependent if there exists a $k$-flat $K$ and $k+2$ distinct integers $i_{1},\dots,i_{k+2}$ such that for all $j\in[k+2]$ we have $K\cap B_{i_{j}{}}\neq\emptyset.$ If a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets is not $k$-dependent then we will call it $k$-independent. A collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ of families of convex sets in $\mathbb{R}^{d}$ is said to satisfy heterochromatic $(\aleph_{0},k+2)$-property if every heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent. Similarly, we will say that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies weak heterochromatic $(\aleph_{0},k+2)$-property if every strictly heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent.
In this paper, we give a number of countably colorful $(\aleph_{0},k+2)$-theorems for $k$-transversals.
Theorem 1.10 (Countably colorful $(\aleph_{0},\cdot)$-theorem).
@@
Let $d\in\mathbb{N}$, and for each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of $(r,R)$-fat closed convex sets in $\mathbb{R}^{d}$.
  • (i)
    For a non-negative integer $k$ with $10\leq k\leq d-$ 1 if $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the heterochromatic $(\aleph_{0},k+2)$-property, then all but finitely many $\mathcal{F}_{n}$’s can be pierced by finitely many $k$-flats.
  • (ii)
    If $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property, then there exist at least two distinct families $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ with $i\neq j$ and each of them can be pierced by finitely many points.
The above theorem is a colorful generalization of Theorem 1.7 by Keller and Perles [KP22].
@@
The following proposition shows that Theorem 1.10 $(ii)$ is tight for families satisfying weak heterochromatic $(\aleph_{0},2)$-property.
Proposition 1.11.
@@
    exactly two distinct $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ are pierceable by finite number of points.
Given Theorem 1.10 $(ii)$ it is natural to ask the following question:
Let $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ be an infinite sequence of families of $(R,r)$-fat closed convex sets satisfying weak heterochromatic $(\aleph_{0},k+2)$-property for $d>k>0$. Does there exist a family $\mathcal{F}_{m}$ from the sequence $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ that has a finite size $k$-transversal?
We show that the above statement is not true for all non-negative $k$ with $11\leq k\leq d-$ 1 even if we restrict ourselves to the case of unit balls.
Theorem 1.12.
@@
Keller and Perles [KP22] showed that fatness assumptions can be removed for families of closed balls and point transversals.
Theorem 1.13 (Keller and Perles [KP22]: $(\aleph_{0},2)$-theorem for closed balls).
Let $\mathcal{F}$ be a family of closed balls in $\mathbb{R}^{d}$ (with no restriction on radii). Then either $\mathcal{F}$ has a point transversal of finite size or it contains an infinite sequence of pairwise disjoint balls.
We will prove a colorful variant of the Theorem 1.13 for families of closed convex sets satisfying heterochromatic $(\aleph_{0},2)$-property. To state the result we will need some additional notations and definitions. For a set $S\subseteq\mathbb{R}^{d}$, the out-radius $R(S)$ is defined as the radius of the smallest closed ball containing $S$, and in-radius $r(S)$ of $S$ is defined as the radius of the largest closed ball contained in $C$. Condition number $\sigma\left(S\right)$ of a set $S$ is defined as $\sigma(S):=\frac{r(S)}{R(S)}$. For a family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$, condition number $\sigma(\mathcal{F})$ of $\mathcal{F}$ is defined as $\sigma(\mathcal{F}):=\min\left\{\sigma(S)\;:\;S\in\mathcal{F}\right\}$.
Theorem 1.14 (Countably colorful $(\aleph_{0},2)$-theorem and bounded condition number).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ satisfying heterochromatic $(\aleph_{0},2)$-property and $\liminf\limits_{n\to\infty}\sigma(\mathcal{F}_{n})>0$. Then all but finitely many $\mathcal{F}_{n}$’s can be pierced by finitely many points.
Corollary 1.15.
Let $\mathcal{F}$ be a family of closed convex sets in $\mathbb{R}^{d}$ with $\sigma(\mathcal{F})>0$. Then either $\mathcal{F}$ has a point transversal of finite size or it contains an infinite sequence of pairwise disjoint convex sets.
Both Theorem 1.10 and Theorem 1.14 require the convex bodies to have nice shape. In Theorem 1.10, we need the convex bodies to be $(r,R)$-fat, and in Theorem 1.14 we require the condition number of the families to be bounded away from zero. It is natural to ask if one can prove a result, similar to the above mentioned theorems, for some natural families of convex bodies that are neither fat nor do they have a lower bound on the condition number. We can show the following heterochromatic $(\aleph_{0},2)$-theorem for axis-parallel boxes.
Theorem 1.16 (Heterochromatic $(\aleph_{0},2)$-theorem for boxes).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of axis-parallel boxes in $\mathbb{R}^{d}$ such that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies heterochromatic $(\aleph_{0},2)$-property. Then all but finitely many $\mathcal{F}_{n}$’s are pierced by finitely many points.
The above result is interesting because axis-parallel boxes are neither fat nor do they have a lower bound on the condition number away from zero. The following impossibility result shows that
  • Theorem 1.16 is not true if boxes are not axis-parallel.
  • Assuming the convex bodies have non-zero volumes is not sufficient to establish a $(\aleph_{0},\cdot)$-theorem.
Theorem 1.17 (Impossibility result).
Remark 1.18.
@@
  • In Section 3 we will prove Theorem 1.10, Proposition 1.11 and Theorem 1.12.
  • Theorem 1.14 and Theorem 1.16 will be proved in Section 4.

### v5

@@ Abstract
An infinite sequence of sets $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence of families $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$, if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. In this paper, we have proved that if for each $n\in\mathbb{N},\mathcal{F}_{n}$ is a family of nicely shaped convex sets in $\mathbb{R}^{d}$ such that each heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ contains at least one $k+2$-tuple that can be pierced by a single $k$-flat ($k$-dimensional affine space) then all but finitely many $\mathcal{F}_{n}$’s can be pierced by finitely many $k$-flats. This result can be considered as a countably colorful generalizations of the $(\aleph_{0},k+2)$-theorem proved by Keller and Perles (Symposium on Computational Geometry 2022). We have also established the tightness of our result by proving a number of no-go theorems.
@@ 1 Introduction
Danzer, Grünbaum and Klee [DGK63] showed that the ”congruent” assumption in Theorem 1.2 can be weakened as long as the sets are of bounded diameter, that is, there exists $R>0$ such that the diameter of all the convex sets in the family is at most $R$.
@@
Observe that for $k$-transversal $q(k)=k+2$ is optimal since any $k+1$ nonempty subset of $\mathbb{R}^{d}$ can be pierced by a single $k$-flat. Keller and Perles proved a similar result for point transversal of families of closed balls (with no restriction on radii).
Theorem 1.8 (Keller and Perles [KP22]: $(\aleph_{0},2)$-Theorem for point transversal).
Suppose $\mathcal{F}$ is a family of closed balls in $\mathbb{R}^{d}$ (with no restriction on the radii) satisfying the $\left(\aleph_{0},2\right)$-property for points. Then $\mathcal{F}$ has a finite-size point transversal.
To complement the above result, Keller and Perles also gave an explicit example showing that the compactness assumption in Theorem 1.7 cannot be weakened. Additionally, Keller and Perles observed that, unlike $(p,q)$-property, for all $10\leq k\leq d-$ 1 there does not exist a constant $c(d,k)>0$ such that if $\mathcal{F}$ is a family of unit balls in $\mathbb{R}^{d}$ satisfying $(\aleph_{0},k+2)$-property with respect to $k$-flats then $\mathcal{F}$ has a $k$-transversal of size at most $c(d,k)$. In an updated version [KP23] Keller and Perles have extended Theorem 1.7 to families of near-balls. A collection $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ is called a family of near-balls if there exists a constant $\alpha>0$ such that for all $C\in\mathcal{F}$, there exist closed balls $B(p_{C},r_{C})$ and $B(p_{C},R_{C})$ satisfying the following two conditions:
$B(p_{C},r_{C})\subseteq C\subseteq B(p_{C},R_{C}),~{}\mbox{and}$ (1)
$R_{C}\leq\min\left\{r_{C}+\alpha,\alpha r_{C}\right\}.$ (2)
Theorem 1.9 (Keller and Perles [KP23]: $(\aleph_{0},k+2)$-Theorem for near balls).
Let $d$ be a natural number and $k$ be a non-negative integer with $10\leq k\leq d-$ 1. Suppose $\mathcal{F}$ be a collection of compact near-balls satisfying the $(\aleph_{0},k+2)$-property for $k$-flats. Then $\mathcal{F}$ has a finite-size $k$-transversal.
Observe that the above result directly gives the following corollary.
Corollary 1.10 (Keller and Perles [KP23]: $(\aleph_{0},k+2)$-Theorem for closed balls).
Let $d$ be a natural number and $k$ be a non-negative integer with $10\leq k\leq d-$ 1. Suppose $\mathcal{F}$ be a collection of closed balls in $\mathbb{R}^{d}$ satisfying the $(\aleph_{0},k+2)$-property for $k$-flats. Then $\mathcal{F}$ has a finite-size $k$-transversal.
Note that Corollary 1.10 is a generalization of Theorem 1.7. Jung and Pálvölgyi [JP23] recently developed a new framework using which one can lift $(p,q)$-type theorems for finite families to prove new $(\aleph_{0},q)$-Theorems for infinite families, and also provided new proofs of already existing results like Corollary 1.10.
Given $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ families of convex sets in $\mathbb{R}^{d}$, we call a collection $B_{1},\dots,B_{p}$ of $p$ convex sets colorful with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ if for all $i\in\left\{1,\dots,p\right\}$ we have $B_{i}\in\mathcal{F}_{i}$. We say $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ satisfies colorful $(p,q)$-property with respect to $k$-flats if every colorful collection of $p$ sets with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ contains $q$ sets that can be pierced by a single $k$-flat. Lovász observed the following interesting variant of Helly’s theorem, called the Colorful Helly’s theorem, in 1974, though the proof first appeared in a paper on Carathéodory’s theorem by Bárány [Bár82].
Theorem 1.11 (Bárány [Bár82]: Colorful Helly’s theorem).
@@
Bárány, Fodor, Montejano, Oliveros, and Pór [BFM+14] proved a colorful $(p,q)$-theorem for families of convex sets in $\mathbb{R}^{d}$.
Theorem 1.12 (Bárány-Fodor-Montejano-Oliveros-Pór [BFM+14]: Colorful $(p,q)$-theorem).
@@
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of compact convex sets in $\mathbb{R}^{d}$. An infinite sequence of compact convex sets $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. A heterochromatic sequence is said to be a strictly heterochromatic sequence if each family has a representative in the sequence, that is, a heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a strictly heterochromatic sequence if for each $n\in\mathbb{N},\;B_{n}\in\mathcal{F}_{n}$.
@@
A sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets in $\mathbb{R}^{d}$ is said to be $k$-dependent if there exists a $k$-flat $K$ and $k+2$ distinct integers $i_{1},\dots,i_{k+2}$ such that for all $j\in[k+2]$ we have $K\cap B_{i_{j}{}}\neq\emptyset.$ If a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets is not $k$-dependent then we will call it $k$-independent. A collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ of families of convex sets in $\mathbb{R}^{d}$ is said to satisfy the heterochromatic $(\aleph_{0},k+2)$-property if every heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent. Similarly, we will say that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},k+2)$-property if every strictly heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent.
We will prove a colorful variant of the Theorem 1.9 for families of closed convex sets satisfying heterochromatic $(\aleph_{0},k+2)$-property with respect to $k$-flats. To state the result we will need some additional notations and definitions. For a set $S\subseteq\mathbb{R}^{d}$, the out-radius $R_{S}$ is defined as the radius of the smallest closed ball containing $S$, and in-radius $r_{S}$ of $S$ is defined as the radius of the largest closed ball contained in $C$. The condition number $\sigma\left(S\right)$ of a set $S$ is defined as $\sigma(S):=\frac{r_{S}}{R_{S}}$. For a family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$, condition number $\sigma(\mathcal{F})$ of $\mathcal{F}$ is defined as the number $\sigma(\mathcal{F}):=\min\left\{\sigma(S)\;:\;S\in\mathcal{F}\right\}$.
Theorem 1.13 (Countably Colorful $(\aleph_{0},k+2)$-Theorem).
@@
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ satisfying the heterochromatic $(\aleph_{0},k+2)$-property and
$\liminf\limits_{n\to\infty}\sigma(\mathcal{F}_{n})>0.$
Then there is a finite set of $k$-flats that pierces all but finitely many $\mathcal{F}_{n}$’s.
Corollary 1.14 (Generalization of Theorem 1.7 and Theorem 1.9).
Let $d$ be a natural number and $k$ be a non-negative integer $k$ with $10\leq k\leq d-$ 1. If $\mathcal{F}$ be a family of closed convex sets in $\mathbb{R}^{d}$ with $\sigma(\mathcal{F})>0$ and $\mathcal{F}$ satisfy $(\aleph_{0},k+2)$-property with respect to $k$-flats then $\mathcal{F}$ has a constant size $k$-transversal.
Using a different argument, Jung and Pálvölgyi [JP23] have given an alternate proof of the above corollary. Note that our proof uses entirely geometric arguments and builds on the approach of Keller and Perles [KP22]. We can also show that the convexity assumption in Theorem 1.13 and Corollary 1.14 cannot be weakened.
Theorem 1.15.
There is an infinite sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of compact connected sets in $\mathbb{R}^{2}$ with condition number $\rho>0$ such that $\{B_{n}\}_{n\in\mathbb{N}}$ satisfies the $(\aleph_{0},2)$ property, but $\{B_{n}\}_{n\in\mathbb{N}}$ is not pierceable by a finite number of points.
Observe that Theorem 1.7, Theorem 1.9, and Theorem 1.13 all require the convex bodies to have nice shape. In Theorem 1.7 and Theorem 1.9, we need the convex bodies to be $(r,R)$-fat and closed balls respectively, and in Theorem 1.13 we require the condition number of the families to be bounded away from zero. It is natural to ask if one can prove a result, similar to the above mentioned results, for some natural families of convex bodies that are neither fat nor do they have a lower bound on the condition number. Chakraborty, Ghosh and Nandi [CGN23a] have proved an $(\aleph_{0},k+2)$-theorem for axis-parallel boxes and axis-parallel $k$-flats. Axis-parallel boxes are neither fat nor do they have a lower bound on the condition number away from zero. The following impossibility result shows that
  • Such a result is not true if the boxes are not axis-parallel.
  • Assuming that the convex bodies have non-zero volumes is not sufficient to establish a $(\aleph_{0},\cdot)$-theorem.
Theorem 1.16 (Impossibility result).
Remark 1.17.
We say that a collection $\mathcal{F}$ of compact convex sets has bounded diameter if there exists an $R>0$ with $\text{diam}(C)<R$ for all $C\in\mathcal{F}$. We say a collection $\mathcal{F}$ of compact convex sets is unbounded if $\bigcup_{C\in\mathcal{F}}C$ is unbounded. As defined in [AGP02], we say $y\in\mathbb{S}^{d-1}$ to be a limiting direction of an unbounded collection $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{d}$ with bounded diameter if there exists an infinite sequence of points $\left\{s_{n}\right\}_{n\in\mathbb{N}}$ in $\mathbb{R}^{d}$ and an infinite sequence of sets $\left\{S_{n}\right\}_{n\in\mathbb{N}}$ from $\mathcal{F}$ satisfying the following conditions:
  • for all $n\in\mathbb{N}$, $S_{n}\in\mathcal{F}$,
  • for all $n\neq m$ in $\mathbb{N}$, $S_{n}\neq S_{m}$,
  • for all $n\in\mathbb{N}$, $s_{n}\in S_{n}$, and
  • $\lim\limits_{n\to\infty}\|s_{n}\|=\infty$ and $\lim\limits_{n\to\infty}\frac{s_{n}}{\|s_{n}\|}=y$
For an unbounded collection $\mathcal{F}$ of compact convex sets with bounded diameter in $\mathbb{R}^{d}$, limiting direction set $LDS(\mathcal{F})$ of $\mathcal{F}$ is defined as
$LDS(\mathcal{F}):=\left\{y\in\mathbb{S}^{d-1}\,:\,y\;\mbox{is a limiting direction of}\;\mathcal{F}\right\}.$
An unbounded collection $\mathcal{F}$ of compact convex sets with bounded diameter is called $k$-unbounded if the vector space spanned by set $LDS(\mathcal{F})$ has dimension at least $k$. As already mentioned earlier in the introduction, Santaló [San40] established the impossibility of getting a Helly Theorem for transversals of arbitrary convex sets with $k$-flats. Aronov, Goodman and Pollack [AGP02] proved the Helly Theorem for $k$-flats for $k$-unbounded families of arbitrary convex compact sets. We show the impossibility of getting $(\aleph_{0},k+2)$-Theorem for $k$-unbounded families of arbitrary convex compact bodies.
Theorem 1.18 (No-go theorem for $k$-unbounded families).
There exists an infinite family $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{3}$ satisfying the following:
  • (a)
    $\mathcal{F}$ is one-unbounded,
  • (b)
    $\mathcal{F}$ satisfies the $(\aleph_{0},3)$-property with respect to lines, and
  • (c)
    $\mathcal{F}$ does not have a finite line transversal.
Given Theorem 1.13 it is natural to ask the following question:
Let $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ be an infinite sequence of families of nicely shaped closed convex sets satisfying
weak heterochromatic $(\aleph_{0},k+2)$-property for $10\leq k\leq d-$ 1. Does there exist a family $\mathcal{F}_{m}$
from the sequence $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ that has a finite size $k$-transversal?
@@
We show the following colorful $(\aleph_{0},2)$-Theorem for families of $(R,r)$-fat convex sets satisfying weak heterochromatic $(\aleph_{0},2)$-property.
Theorem 1.19 (Strongly Colorful $(\aleph_{0},2)$-Theorem).
Let $d\in\mathbb{N}$, and for each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of $(r,R)$-fat closed convex sets in $\mathbb{R}^{d}$. If $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property then there exist at least two distinct families $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ with $i\neq j$ and each of them can be pierced by finitely many points.
The following proposition shows that Theorem 1.19 is tight for families satisfying weak heterochromatic $(\aleph_{0},2)$-property.
Proposition 1.20 (Tightness of Strongly Colorful Theorem 1.19).
@@
    exactly two distinct families $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ are pierceable by finite number of points, and the rest of the families from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ are not pierceable by a finite number of points.
We show the impossibility of getting an analog of Theorem 1.19 for all natural numbers $k$ satisfying $0<k<d$ even if we restrict ourselves to families of unit balls.
Theorem 1.21 (No-go result for strongly colorful $(\aleph_{0},k+2)$-Theorem).
@@
  • In Section 3 we will prove of Theorem 1.13.
  • In Section 4, we prove Theorem 1.19 and Proposition 1.20.
  • We will prove Theorem 1.15, Theorem 1.16, Theorem 1.18 and Theorem 1.21 in Section 5.